OpenStudy (nurali):

1. A simple random sample of 400 students is taken at a large university. The average height of the sampled students is 68 inches and the SD is 2 inches. The distribution of heights in the sample follows the normal curve very closely. An approximate 68% confidence for the average height of students at the university is 68 inches plus or minus ____________ inches. Approximately 68% of the students in the sample have heights in the range 68 inches plus or minus ____________ inches.

OpenStudy (kropot72):

The 68% confidence interval for the mean is \[68-0.997 \frac{2}{\sqrt{400}}< \mu <68+0.997 \frac{2}{\sqrt{400}}\]

OpenStudy (anonymous):

An approximate 68% confidence for the average height of students at the university is 68 inches plus or minus 0.0997 inches. Approximately 68% of the students in the sample have heights in the range 68 inches plus or minus 2 inches.